Margin of safety

Organic Health Care Products Inc. expects to maintain the same inventories at the end of 20Y8 as at the beginning of the year. The total of all production costs for the year is therefore assumed to be equal to the cost of goods sold. With this in mind, the various department heads were asked to submit estimates of the costs for their departments during 20Y8. A summary report of these estimates is as follows:

Estimated Fixed Cost Estimated Variable Cost
(per unit sold)
Production costs:
Direct materials $ 8.00
Direct labor 3.00
Factory overhead $ 200,000 1.50
Selling expenses:
Advertising 1,450,000
Sales salaries and commissions 93,000 1.85
Travel 340,000
Miscellaneous selling expense 2,000 0.10
Administrative expenses:
Office and officers' salaries 300,000
Supplies 10,000 0.50
Miscellaneous administrative expense 5,000 0.05
Total $2,400,000 $15.00

It is expected that 400,000 units will be sold at a price of $25 a unit. Maximum sales within the relevant range are 500,000 units.

Determine the following based upon the estimates for 20Y8:

1. Margin of safety for 20Y8 in units sold. $
2. Margin of safety for 20Y8 in sales dollars. $
3. Margin of safety for 20Y8 as a percentage of sales. %
Answer #1

1) Break even unit = Fixed cost/Contribution margin per unit = 2400000/(25-15) = 240000 Units

Margin of safety in units = 400000-240000 = 160000 Units

2) Margin of safety in dollars = 160000*25 = $4000000

3) Margin of safety (%) = 160000/400000 = 40%